the work done when 6.5 g of zinc reacts with dilut HCl in an open beaker at 298k is

Grade:11

1 Answers

Vikas TU
7 years ago
moles of zn reacted = > 6.5/65 = 0.1 mol.
from ideal gas eqn.
PV = nRT
1*V = 0.1*0.0821*298
V = 2.46 litre.
Work Done = -Pext.*V => -1*2.46 = -2.46 l/atm
or
W = -2.46 * 100 => -246 Joules (approx.)